# Shuffle Mamba: State Space Models with Random Shuffle for Multi-Modal Image Fusion

### v1

@@ Introduction
The quantitative comparison results of four metrics on the MIF dataset are shown in Table 3. The proposed method performs well across almost all metrics, demonstrating its effectiveness in medical image fusion. In our experimental results, a higher VIF indicates closer alignment with human perception. Improved SCD, Qabf, and SSIM scores suggest that the fused image maintains higher similarity and experiences less distortion compared to the original images. Figure 6 presents a qualitative comparison of several methods on the MRI-PET dataset. Our method exhibits superior visual quality, a finding that is corroborated by the experimental metrics.
Ablation Experiments
To verify the effectiveness of the Random Shuffle operation, we conducted corresponding ablation experiments. The core operations of these experiments involved removing the shuffle operations from the three main modules. Notably, in the experiment where RM-related components were removed, the RM block in the RCIM module was excluded. The experimental results are presented in Table 4. Removing the shuffle operation resulted in a significant decrease in model performance, indicating that the new scanning strategy designed under the Shuffle Mamba framework effectively enhances the quality of multi-modal image fusion.
Due to Monte Carlo averaging, Shuffle Mamba makes trading memory and computational resources possible for improved performance. We studied the relationship between the number of samples, the PSNR index, and resource consumption. We experimented five times for each sample size, calculating the average and standard deviation of diverse metrics. Figure 7 (a) illustrates the PSNR trend, while Figure 7 (b) and (c) details the corresponding memory usage and the time required to process each image. As the number of samples increases, performance and resource consumption rise, enabling a trade-off between performance and efficiency. Additionally, Monte Carlo averaging enhances the theoretical robustness of the random shuffle operation, effectively improving the mean of PSNR while reducing the variance of the fusion result. More experimental results can be found in the supplementary materials.
@@
Conclusion
In this paper, we replace the fixed scanning strategy used in current Mamba-based methods with a random shuffle-based scanning method and design a new Shuffle Mamba framework. This approach mitigates the bias introduced by fixed scanning methods and provides an global receptive field for multi-modal image fusion. During testing, we employ Monte Carlo averaging to account for the introduced random factors. Extensive experiments on two tasks demonstrate that our approach outperforms state-of-the-art methods and exhibits strong generalization capabilities.

### v2

@@ IV-C2 Medical Image Fusion
The quantitative comparison results for four metrics on the MIF dataset are presented in Table IV. The proposed method demonstrates strong performance across nearly all metrics, highlighting its effectiveness in medical image fusion. In our experimental results, a higher VIF score indicates closer alignment with human visual perception. Enhanced SCD, Qabf, and SSIM scores suggest that the fused image preserves greater similarity and undergoes less distortion compared to the original images. Qualitative comparisons of several methods on the MRI datasets are shown in Figure 9, Figure 10, and Figure 11. Our method consistently delivers superior visual quality, a finding further supported by the experimental metrics. Furthermore, to evaluate perceptual quality, we performed a user study involving ten participants with medical backgrounds, including graduate students and clinical researchers. Each of the 72 MIF test pairs (ours vs. CDDFuse) was randomly displayed for pairwise comparison. The proposed method was preferred in 60 out of 72 cases (83.3%), primarily due to clearer anatomical boundaries and improved soft-tissue visibility, supporting the quantitative superiority with perceptual evidence.
IV-C3 Infrared and Visible Image Fusion
To further evaluate the generalization capability of the proposed method across different multi-modal image fusion tasks, we additionally conducted experiments on Infrared and Visible Image Fusion (IVIF). In the IVIF setting, all methods were trained on the MSRS dataset [49] and tested on three benchmark datasets: MSRS, RoadScene [61], and M3FD [31]. Besides the comparison methods used in the MIF experiments, including U2Fusion [60], SwinFusion [36], CDDFuse [71], PSLPT [53], FusionMamba [43], and TC-MoA [77], we further extended the comparison to include SemLA [59], MLFuse [21], and Up-Fusion [23].
Table V presents the quantitative results on three datasets. It can be observed that the proposed method consistently achieves the best overall performance across all evaluation metrics and datasets. Specifically, on the MSRS dataset, our method attains the highest scores in SCD, VIF, Qabf, and SSIM, indicating superior capability in preserving complementary modal information and structural consistency. For the RoadScene dataset, our method ranks first on three metrics and achieves the second-best performance on Qabf, demonstrating robust generalization to complex driving scenarios with varying illumination and scene dynamics. These results validate that the proposed framework not only excels in the pan-sharpening and MIF but also exhibits strong cross-modal generalization and adaptability in the IVIF task.
IV-D Ablation Experiments
To evaluate the effectiveness of the Random Shuffle mechanism, we conducted ablation experiments by removing the shuffle operation from the RM, RCIM, and RMIM modules, individually and in combination. As shown in Table VI, removing the shuffle operation consistently leads to performance degradation across all three datasets, confirming that the proposed scanning strategy is a key contributor to fusion quality. Notably, the degree of performance drop varies across datasets and modules. On WorldView-II, removing the shuffle in the RCIM module results in the largest decline. When multiple shuffle operations are removed simultaneously, the performance loss becomes more pronounced, demonstrating that these modules provide complementary gains. These observations confirm that the random-shuffle scanning strategy enhances global feature diversity and effectively reduces modality bias in different fusion stages.
In addition, we further evaluated the influence of different scanning strategies through a comparative ablation study, as presented in Table VII. In this experiment, the default Random Shuffle Scanning (RSS) was replaced with Sequential Scanning (SS), Bidirectional Scanning (BS), and Diagonal Scanning (DS), respectively. The results show that RSS consistently outperforms all deterministic scanning strategies in all metrics. Specifically, while deterministic strategies such as BS and DS yield marginal improvements over SS, none of them can match the performance achieved by RSS. This confirms that introducing stochasticity through the Random Shuffle mechanism enables more diverse and global spatial interactions, effectively improves global context aggregation, and ultimately leads to superior fusion performance.
Figure 12: The effective receptive fields of different scanning schemes.
To validate the effectiveness of the proposed random shuffle scanning strategy, we provide a comparative visualization of the effective receptive fields (ERFs) under different scanning schemes. As illustrated in Figure 12, the long-range dependencies captured by 2D scanning are primarily concentrated along horizontal and vertical directions. At the same time, diagonal and zigzag scans exhibit global receptive fields aligned with their respective diagonal paths. In contrast, our method produces a more uniformly distributed ERF, characterized by overall darker responses without prominent directional preferences. This demonstrates that the proposed random shuffle strategy effectively mitigates the local bias inherent in fixed scanning patterns, facilitating more balanced global modeling.
@@ IV-D Ablation Experiments
Due to Monte Carlo averaging, Shuffle Mamba makes trading memory and computational resources possible for improved performance. We studied the relationship between the number of samples, the PSNR index, and resource consumption. We conducted five experiments for each sample size, calculating the average and standard deviation of various metrics. Figure 13 (a) illustrates the PSNR trend, while Figure 13 (b) and (c) detail the corresponding memory usage and the time required to process each image. As the number of samples increases, performance and resource consumption rise, enabling a trade-off between performance and efficiency. Additionally, Monte Carlo averaging enhances the theoretical robustness of the random shuffle operation, effectively improving the mean of PSNR while reducing the variance of the fusion result. It is noteworthy that even when the number of Monte Carlo samples is set to 1, the network still performs well. This is because the overall architecture consists of multiple stacked modules, each with independent random shuffle scans. Therefore, multiple instances of random shuffling and recovery occur within the network, which further ensures stable performance and demonstrates the robustness of the proposed method.
V Conclusion
In this paper, we replace the fixed scanning strategy used in current Mamba-based methods with a random shuffle-based scanning method and design a new Shuffle Mamba framework. This approach mitigates the bias introduced by fixed scanning methods and provides a global receptive field for multi-modal image fusion. During testing, we employ Monte Carlo averaging to account for the random factors introduced. Extensive experiments on two tasks demonstrate that our approach outperforms state-of-the-art methods and exhibits strong generalization capabilities.
Despite these advantages, our method has certain limitations. The primary drawback lies in the computational cost associated with the Monte-Carlo averaging during inference. As analyzed in Figure 13, while increasing the number of scan samples improves the results and reduces variance, it inevitably increases inference time and memory consumption linearly. Achieving the theoretical upper bound of performance requires multiple forward passes, which may limit real-time applications on resource-constrained edge devices. In future work, we aim to explore more efficient scanning strategies that maintain unbiased global perception without requiring repetitive sampling. Additionally, inspired by recent trends, we plan to extend the Shuffle Mamba framework to handle complex fusion scenarios, such as misaligned inputs and extreme weather conditions.